Circular No. 5864

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRS 1009–45 I. Lapshov, S. Sazonov, and R. Sunyaev, on behalf of the GRANAT

WATCH team (Space Research Institute, Moscow; and Danish Space Research Institute, Lyngby), report: “A new hard x-ray transient source GRS 1009–45 (α = 10^{h}09^{m}, δ = −44^{o}55′, equinox 1950.0; radius of uncertainty 1^{o}) was detected by the GRANAT/WATCH All Sky Monitor during observations performed on Sept. 12-13. The source’s intensity in the band 8-20 keV (in Crab units) is as follows: Sept. 12-13 UT, 0.6 ± 0.1; 13-14, 0.7 ± 0.1; 15.1, 0.6 ± 0.1. Similarly, in the 20-60 keV band: Sept. 12.0, 0.4 ± 0.1; 12-14, 0.8 ± 0.15; 15.1, 0.7 ± 0.1. The source was too weak on Sept. 11 to be detected by WATCH. The spectrum of the source is harder than spectra of Crab Nebula and Vela X-1 observed simultaneously with the same detector; its hardness is comparable to that of Cygnus X-1.”

B. A. Harmon, Marshall Space Flight Center (MSFC), NASA; S. N. Zhang, Universities Space Research Association and MSFC; G. J. Fishman, MSFC; and W. S. Paciesas, University of Alabama at Huntsville and MSFC, report for the Compton Gamma Ray Observatory BATSE team: “We have detected strong hard x-ray/γ-ray emission from GRS 1009–45 in the range 20 to 500 keV. The source was located using the newly-developed BATSE occultation transform imaging system, and was determined to be centered at α = 10^{h}14^{m}, δ = −45^{o}.5 (equinox 2000.0), which is within the WATCH error circle. The BATSE location error is ± 0^{o}.2 in the northwest-southeast direction and ± 0^{o}.4 in the perpendicular direction. Daily flux averages, relative to the Crab nebula, are (errors ∼ 0.1 Crab): Sept. 12, 0.5; 13, 0.7; 14, 0.9; 15, 0.9; 16, 0.8, with power-law indices ranging from –2.2 to –2.6. The Sept. 13-16 average spectrum is well fit by a single power law with index –2.5 to at least 500 keV.”

R CORONAE BOREALIS Further visual magnitude estimates (cf. IAUC 5851): Aug. 31.83 UT,

7.0 (L. Szentasko, Budapest, Hungary); Sept. 4.82, 7.2 (H. Bengtsson, Hisings Backa, Sweden); 9.83, 7.4 (A. Pereira, Cabo da Roca, Portugal); 11.20, 7.4 (C. E. Spratt, Victoria, BC); 13.86, 7.1 (Pereira); 16.19, 7.2 (Spratt); 18.19, 6.9 (Spratt).

1993 September 18 Daniel W. E. Green